OpenStudy (anonymous):

A 3/8 foot thick board was stacked on top of a 1/3 foot thick board. What was the total thickness of the stack?

OpenStudy (anonymous):

first find a common denominator to add the fractions 3/8 = 9/24 1/3 = 8/24 so what is \[\frac{ 9 }{ 24 }+\frac{ 8 }{ 24 } = ?\]